# Next-Gen Software Engineering: AI-Assisted Big Models

### v1

@@ 5 AI-Assisted Big Models in Software Engineering
Hence, as AI is increasingly applied to Big Models, it is necessary to incorporate the specific aspects of modelling into the AI4SE taxonomy, which currently does not distinguish between modelling and other tasks related to software development. At this early stage of development, it is not yet clear whether modelling should be treated as a separate task or like software quality and software security as a cross-cutting concern affecting numerous facets within the AI4SE taxonomy.
@@ 6 Summary
After reviewing the state of the practice of model-driven software engineering, big code on open source software platforms, and the application of AI in software engineering, the taxonomy of AI-assisted software engineering is developed and used to categorise recent publications. In addition, the concept of Big Models is defined and explored with a view to future opportunities for further adoption of model-driven software engineering. Finally, recent research on the application of AI approaches to big models is reviewed and categorised.
Future work will further explore the evolution of Big Models and AI applications in MDSE, potentially leading to an update of the new AI4SE taxonomy.
Acknowledgements
The ideas presented in this paper have been developed through constructive dialogue in the Feldafinger Kreis, the German Testing Board and the Association for Software Quality and Education. The author acknowledges that while authored by her, the writing process was aided by AI tools, specifically ResearchRabbit for determining related work, and ChatGPT and DeepL for fine-tuning the wording.
The author has no competing interests to declare that are relevant to the content of this article.

### v2

@@ 5 AI-Assisted Model-Driven Software Engineering
And as AI is increasingly applied to Big Models, it remains to be seen whether it becomes necessary to include more specifics of modelling in the AI4SE taxonomy, which currently considers modelling as an optional facet in software development, operations and SE processes. At this early stage of development, it is not clear whether modelling needs to be a separate facet or can remain a cross-cutting concern affecting numerous facets within the AI4SE taxonomy.
6 Related Work
The growing intersection of artificial intelligence (AI) and software engineering (SE) represents a highly active field in research and tool development. The sheer volume of research results has necessitated the creation of surveys [2, 86, 30, 3, 31, 65, 32]that review, sort and structure the state of research in this field. Natural language processing (NLP) techniques, as well as (semi-)formal language processing on large code bases (Big Code) for AI-augmented software engineering, are used to address areas such as code generation, defect detection, code repair, code documentation, software testing, software quality and the challenges related to software security and piracy. Another focus is the application of AI in MDSE [18, 26], examining their use in automating tasks like model generation, transformation, and management, alongside considerations for integration and potential challenges like hallucinations.
It is widely acknowledged that artificial intelligence, and machine learning in particular, is becoming increasingly significant in the field of software engineering. The papers emphasise the potential of AI to automate tasks, enhance code quality, and support developers. It is further noted that AI has the potential to relieve developers from repetitive work, thereby allowing them to focus on higher-level design and problem-solving. Additionally, there is a discussion of the potential of LLMs for code completion, code generation, and other tasks, as well as the challenges and shortcomings of current LLMs. A recurring theme in the literature is the call for more empirical data to support the advancement of this field. Hence, the necessity for a data-driven approach to SE is emphasised by several papers. Also, some papers express concerns about potential biases in AI systems that could have an impact on SE and the resulting software quality.
The papers also exhibit different foci: due to the limited dissemination of MDSE, the majority focusses on the use of AI directly on code, using extensive code bases to train models for code generation and analysis. A smaller portion analyses the application of AI within the context of MDSE, exploring how AI can automate model transformations and facilitate model creation. Furthermore, different taxonomies and classifications are discussed: The papers propose different ways to classify the application of AI in SE, which demonstrates that the field is developing and still lacks a unified framework where AI4SE aims at becoming it. Additionally, there is a divergence of opinion regarding the desired level of automation. Some papers advocate for the utilisation of AI as a tool to assist developers, while others envision the development of fully autonomous systems capable of making decisions without human intervention. Generally, the papers are focused on automating repetitive tasks in software development and reducing errors that may arise from human actions.
More specifically, [32] was one of the first to bridge the gap between AI and SE arguing for their fusion to meet emerging software needs. It explores the potential of integrating AI and SE techniques and tools to create more sophisticated and user-friendly software. [65] discusses how the traditionally separate disciplines of AI and SE are increasingly intertwined. It examines the use of AI techniques to improve SE processes (e.g., project management) and the use of SE methods to develop and manage AI systems (e.g., agent-oriented software engineering). [3] introduces the "naturalness hypothesis" (see section 4) and uses it to justify a data-driven approach to SE. It presents a taxonomy of probabilistic models (code generation, representation, and pattern mining models) and describes their underlying assumptions and applications to SE tasks like code completion, bug detection, and program synthesis. The paper identifies key challenges, such as data sparsity and the need to bridge symbolic programming language representations with ML representations. [86] updates the use of the "naturalness hypothesis" for code generation, completion, translation, refinement, summarisation, defect and clone detection by GitHub Copilot and AlphaCode. It also discusses the challenges and opportunities, including computational cost, quality measurement, software security, piracy concerns, and integration with existing development tools such as Xcode. [2] examines current AI applications across the software development lifecycle, from requirements analysis to maintenance, using tools such as GitHub Copilot and AlphaCode. It analyses the benefits (such as increased productivity and improved code quality) and challenges (such as data quality, ethical concerns and skills gaps) of AI adoption in SE. It also aims to provide a holistic framework for integrating AI responsibly, balancing automation with human oversight, and offers actionable recommendations for practitioners.
Wrt. MDSE, [18] explores the history and future of automation within MDSE. It focuses on the transformative potential of AI for MDSE, addresses challenges such as data scarcity and human-computer interaction, and aims to stimulate further research into the automation capabilities of MDSE through AI and extending its application to different domains and new computing paradigms. [26] examines the burgeoning intersection of LLMs and MDSE, focusing on the automation of tasks such as model repository classification and recommender system development. It explores how LLMs can improve the management, exploration and evolution of entire modelling ecosystems, thereby transforming software engineering practices.
As far as taxonomies for SE and AI are concerned, only two have been made explicit: [31] presents the AI-SEAL taxonomy based on three key facets: point of application (process, product, or runtime), five types of AI technology used (e.g. AI tribes such as symbolist or connectionist), and ten levels of automation (e.g. from "human considers alternatives, makes and implements decision" to "computer makes and implements decision if it feels it should, and informs human only if it feels this is warranted"). In AI4SE, we expand the point/target of the application by detailing the activities involved, we expand the types of AI by adopting a more natural classification to include, for example, agentic AI, and we limit the levels of automation to the more established levels of autonomy when discussing AI applications. In addition, we add the purpose of the AI application, which guides the design, configuration and evaluation of an AI application in SE.
[30] explores the current and future potential of software development bots ("DevBots") as a form of agentic AI. It classifies DevBots according to their purpose, initiation, communication methods, and level of intelligence. It contrasts current DevBots with a vision of an "ideal" DevBot, which would be autonomous, adaptive, and possess both technical and social skills, acting more like artificial team members than mere tools. Like AI4SE, the DevBots taxonomy uses a facet approach to classification, but is restricted to bots, which limits its scope. It also uses a facet for purpose (although this is different from the purpose in AI4SE) and a facet for intelligence (which is limited to only two levels of autonomy/automation). It has two other bot-specific facets for bot initiation and communication.
@@ 7 Summary
It is anticipated that the integration of AI will enhance the accessibility of software development to a more heterogeneous user base. By automating tasks, AI can free up developers to focus on higher-level design and problem-solving, leading to a more enjoyable and creative development process. Furthermore, it is regarded as the road towards AI-augmented MDSE bridging the gap between higher-level models and concrete code. However, it is crucial to acknowledge the necessity of addressing the ethical implications of AI in SE, particularly in the context of bias, fairness, transparency, and privacy when developing and deploying AI-augmented software tools. The utilisation of pair programming and/or pair modelling with AI has been identified as a catalyst for enhanced collaborative endeavours between humans and AI. In order to facilitate a more extensive integration of AI into tools, a number of challenges must be surmounted. These include the establishment of interoperability among tools, the creation of standards, and the provision of user-friendly coding and modelling tools. In the future, AI-augmented software lifecycle frameworks will (1) have repetitive and often time-consuming tasks being automated, (2) provide intelligent decision-making by analysing Big Code and/or Big Models for deeper insights, (3) provide enhanced quality of software products by detecting requirements inconsistencies. design flaws, software defects, or other issues, early in the development lifecycle, (4) facilitate management of the increasing complexity of modern systems by providing analysis, monitoring and optimisation tools, and (5) enable continuous improvement by feedback loops and adaptive learning, fostering iterative improvement in software products and software processes.
Following a review of the current state of model-driven software engineering practice, this paper examines big code on open-source software platforms and the application of AI in software engineering.The novel taxonomy of AI-assisted software engineering, AI4SE, is developed and used to categorise highly cited and recent publications. Furthermore, the concept of Big Models is defined and explored with a view to future opportunities for further adoption of model-driven software engineering, taking advantage of AI gains. Finally, selected research on AI approaches for MDSE is reviewed and categorised using the AI4SE taxonomy.
The novel AI4SE taxonomy aims to classify and understand the different ways in which AI is used in SE, recognising the growing overlap between AI and SE. Future work will focus on the development of Big Models and AI applications in MDSE and/or detail the other activities in SE.
Acknowledgement
The ideas presented in this paper have been developed through constructive dialogue in the Feldafinger Kreis, the German Testing Board and the Association for Software Quality and Education. The author acknowledges that while authored by her, the writing process was aided by (AI) tools, specifically Google Scholar and ResearchRabbit for determining related work, and NotebookLM, ChatGPT and DeepL for fine-tuning the wording. The author has no competing interests to declare that are relevant to the content of this article.